# Finite-Time Analysis of Simultaneous Double Q-learning

### v1

@@ C.3 Convergence of the lower comparison system
$\rho^{k-1}\biggl{(}\frac{(k-1)k(2k-1)}{6}\biggl{)}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$+\rho^{k-1}\biggl{(}\frac{(k-1)k}{2}\biggl{)}\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$+\frac{2\rho^{k}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}\leq 3\biggl{(}\frac{\rho^{k-3}2k^{3}4\lvert S\times A\rvert^{3/2}}{(1-\gamma)}\biggl{)}$
@@
$\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]$ $\leq\frac{16\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{24\rho^{k-3}k^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}$ (32)
@@ C.4 Convergence of the upper comparison system
$\quad-\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B}),\quad Q_{0}^{A_{U}}-Q_{0}^{A_{L}}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$
@@
$\quad-\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q^{A}})(Q_{k}^{A}-Q_{k}^{B}),\quad Q_{0}^{B_{U}}-Q_{0}^{B_{L}}\in\mathbb{R}^{\lvert S\rvert\lvert A\rvert},$ (33)
@@
$\mathbb{E}[\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}]$ $\leq\frac{40\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{1-\gamma}.$ (34)
@@
$=(I-\alpha D)(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})+\alpha\gamma DP\{\Pi_{Q_{k}^{B}}(Q_{k}^{A_{U}}-Q^{*})+\Pi_{Q_{k}^{B}}(Q_{k}^{A_{L}}-Q^{*})$
$\quad-\Pi_{Q_{k}^{B}}(Q_{k}^{A_{L}}-Q^{*})-\Pi_{Q^{*}}(Q_{k}^{A_{L}}-Q^{*})\}-\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$
$=(I-\alpha D)(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})+\alpha\gamma DP\Pi_{Q_{k}^{B}}(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})$ (35)
$\quad+\alpha\gamma DP(Q_{k}^{A_{L}}-Q^{*})(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})-\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$
@@
$\mathbb{E}[\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}]$ $\leq\frac{8\gamma^{3}d_{\text{max}}^{3}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$\quad+\rho^{k-1}k\frac{2\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}\alpha}{(1-\gamma)}$
$\quad+\rho^{k-1}\frac{(k-1)^{2}k(k-2)}{2}\frac{1}{6}\frac{4\rho^{-3}\alpha^{3}\gamma d_{\text{max}}^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\rho^{k-1}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{1-\gamma}\frac{1}{2}\frac{k(k-1)(k-2)}{3}$
$\quad+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}$
$\quad+\rho^{k-1}\biggl{(}\frac{(k-1)k(2k-1)}{6}\biggl{)}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\rho^{k-1}\biggl{(}\frac{(k-1)k}{2}\biggl{)}\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$ (38)
@@
$\frac{8\gamma^{3}d_{\text{max}}^{3}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{4\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}\leq 5\biggl{(}\frac{8\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}\biggl{)}$
@@
$\rho^{k-1}k\frac{2\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}\alpha}{(1-\gamma)}+\rho^{k-1}\frac{(k-1)^{2}k(k-2)}{2}\frac{1}{6}\frac{4\rho^{-3}\alpha^{3}\gamma d_{\text{max}}^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$\quad+\rho^{k-1}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{1-\gamma}\frac{1}{2}\frac{k(k-1)(k-2)}{3}$
$+\rho^{k-1}\biggl{(}\frac{(k-1)k(2k-1)}{6}\biggl{)}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$
$+\rho^{k-1}\biggl{(}\frac{(k-1)k}{2}\biggl{)}\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}\leq 5\biggl{(}\frac{\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}4}{1-\gamma}\biggl{)}$
@@
$\mathbb{E}[\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}]$ $\leq\frac{40\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{1-\gamma}$ (39)
@@
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{16\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{24\rho^{k-3}k^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}$
$\quad+\frac{40\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{1-\gamma}$ (40)
@@
$\frac{16\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{40\gamma d_{\text{max}}\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{4\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}\leq 3\biggl{(}\frac{40\lvert S\times A\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}\biggl{)}$
@@
$\frac{24\rho^{k-3}k^{3}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert S\times A\rvert^{3/2}}{1-\gamma}\leq 2\biggl{(}\frac{24\rho^{k-4}k^{4}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}\biggl{)}$
@@
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{120\alpha^{1/2}\lvert S\times A\rvert}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{48\rho^{k-4}k^{4}\lvert S\times A\rvert^{3/2}}{(1-\gamma)}$

### v2

@@ C.3 Convergence of the lower comparison system
$\rho^{k-1}\biggl(\frac{(k-1)k(2k-1)}{6}\biggl)\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}+\rho^{k-1}\biggl(\frac{(k-1)k}{2}\biggl)\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}$
$+\frac{2\rho^{k}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}\leq 3\biggl(\frac{\rho^{k-3}2k^{3}4\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}\biggl)$
@@
$\mathbb{E}[\lVert Q_{k}^{A_{L}}-Q^{*}\rVert_{\infty}]$ $\leq\frac{16\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{24\rho^{k-3}k^{3}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}.$ (33)
@@ C.4 Convergence of the upper comparison system
$\quad-\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B}),\quad Q_{0}^{A_{U}}-Q_{0}^{A_{L}}\in\mathbb{R}^{\lvert\mathcal{S}\rvert\lvert\mathcal{A}\rvert}$,
@@
$\quad-\alpha\gamma DP(\Pi_{Q_{k}^{*}}-\Pi_{Q^{A}})(Q_{k}^{A}-Q_{k}^{B}),\quad Q_{0}^{B_{U}}-Q_{0}^{B_{L}}\in\mathbb{R}^{\lvert\mathcal{S}\rvert\lvert\mathcal{A}\rvert}$, (34)
@@
$\mathbb{E}[\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}]$ $\leq\frac{40\gamma d_{\mathrm{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\mathrm{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\mathrm{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}.$ (35)
@@
$Q_{k+1}^{A_{U}}-Q_{k+1}^{A_{L}}$ $=(I-\alpha D)(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})+\alpha\gamma DP\{\Pi_{Q_{k}^{B}}(Q_{k}^{A_{U}}-Q^{*})$
$\quad-\Pi_{Q^{*}}(Q_{k}^{A_{L}}-Q^{*})\}-\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$
$=(I-\alpha D)(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})+\alpha\gamma DP\{\Pi_{Q_{k}^{B}}(Q_{k}^{A_{U}}-Q^{*})$
$\quad+\Pi_{Q_{k}^{B}}(Q_{k}^{A_{L}}-Q^{*})-\Pi_{Q_{k}^{B}}(Q_{k}^{A_{L}}-Q^{*})-\Pi_{Q^{*}}(Q_{k}^{A_{L}}-Q^{*})\}$
@@
$=(I-\alpha D)(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})+\alpha\gamma DP\Pi_{Q_{k}^{B}}(Q_{k}^{A_{U}}-Q_{k}^{A_{L}})+\alpha\gamma DP(Q_{k}^{A_{L}}-Q^{*})(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})$
$\quad-\alpha\gamma DP(\Pi_{Q_{k}^{B}}-\Pi_{Q^{*}})(Q_{k}^{A}-Q_{k}^{B})$
@@
$\mathbb{E}[\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}]$ $\leq\frac{8\gamma^{3}d_{\text{max}}^{3}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}$
$\quad+\rho^{k-1}k\frac{2\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}\alpha}{(1-\gamma)}+\rho^{k-1}\frac{(k-1)^{2}k(k-2)}{2}\frac{1}{6}\frac{4\rho^{-3}\alpha^{3}\gamma d_{\text{max}}^{3}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}$
$\quad+\rho^{k-1}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}\frac{1}{2}\frac{k(k-1)(k-2)}{3}$
$\quad+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{8\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}$
$\quad+\rho^{k-1}\biggl(\frac{(k-1)k(2k-1)}{6}\biggl)\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}$
$\quad+\rho^{k-1}\biggl(\frac{(k-1)k}{2}\biggl)\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}$ (38)
@@
$\frac{8\gamma^{3}d_{\text{max}}^{3}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{4\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{3/2}(1-\gamma)^{5/2}}+\frac{8\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}$
$\quad+\frac{8\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{5/2}(1-\gamma)^{7/2}}$
$\leq 5\biggl(\frac{8\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}\biggr).$
@@
$\rho^{k-1}k\frac{2\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}\alpha}{1-\gamma}+\rho^{k-1}\frac{(k-1)^{2}k(k-2)}{12}\frac{4\rho^{-3}\alpha^{3}\gamma d_{\text{max}}^{3}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}$
$\quad+\rho^{k-1}\frac{k(k-1)(k-2)}{6}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}+\rho^{k-1}\frac{(k-1)k(2k-1)}{6}\frac{4\rho^{-2}\alpha^{2}\gamma^{2}d_{\text{max}}^{2}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}$
$\quad+\rho^{k-1}\frac{(k-1)k}{2}\frac{4\rho^{-1}\alpha\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}$
$\leq 5\biggl(\frac{4\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}\biggr).$
@@
$\mathbb{E}[\lVert Q_{k}^{A_{U}}-Q_{k}^{A_{L}}\rVert_{\infty}]$ $\leq\frac{40\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}.$ (39)
@@
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{16\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{7/2}(1-\gamma)^{9/2}}+\frac{24\rho^{k-3}k^{3}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}+\frac{4\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{1/2}(1-\gamma)^{3/2}}$
$\quad+\frac{40\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert\alpha^{1/2}}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}$ (40)
@@
$\frac{16\gamma\,d_{\max}\lvert\mathcal{S}\times\mathcal{A}\rvert\,\alpha^{1/2}}{d_{\min}^{7/2}(1-\gamma)^{9/2}}+\frac{40\gamma\,d_{\max}\lvert\mathcal{S}\times\mathcal{A}\rvert\,\alpha^{1/2}}{d_{\min}^{9/2}(1-\gamma)^{11/2}}+\;\frac{4\,\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\min}^{1/2}(1-\gamma)^{3/2}}\;\leq\;3\biggl(\frac{40\,\lvert\mathcal{S}\times\mathcal{A}\rvert\,\alpha^{1/2}}{d_{\min}^{9/2}(1-\gamma)^{11/2}}\biggr).$
@@
$\frac{24\rho^{k-3}k^{3}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}+\frac{20\rho^{k-4}k^{4}\alpha\gamma d_{\text{max}}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{1-\gamma}\leq 2\biggl(\frac{24\rho^{k-4}k^{4}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}\biggl).$
@@
$\mathbb{E}[\lVert Q_{k}^{A}-Q^{*}\rVert_{\infty}]$ $\leq\frac{120\alpha^{1/2}\lvert\mathcal{S}\times\mathcal{A}\rvert}{d_{\text{min}}^{9/2}(1-\gamma)^{11/2}}+\frac{48\rho^{k-4}k^{4}\lvert\mathcal{S}\times\mathcal{A}\rvert^{3/2}}{(1-\gamma)}.$